OpenStudy (anonymous):

I have an Economics question but I want to ask it here since more people come to this group...

OpenStudy (anonymous):

To calculate the equilibrium GDP given the following information: C=2,500+0.80Y I=2,225 G=2,000 NX=-775 I would just add them up and when I get 0.80Y + 5,950, I would set it to zero, subtract 5,950 from both sides and divide by 0.80, right?

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@JuanitaM

OpenStudy (anonymous):

@Mertsj

OpenStudy (bahrom7893):

Post the equation for GDP. The data's useless without the equation.

OpenStudy (anonymous):

C+I+G+NX

OpenStudy (anonymous):

is what im getting browsing the internet

OpenStudy (bahrom7893):

Then just add them

OpenStudy (anonymous):

C=2,500+0.80Y There is a Y ^^^^ I was asking about that.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@mathg8 @saifoo.khan

OpenStudy (anonymous):

im gonna post the question like it is on paper

OpenStudy (anonymous):

i need help with #10 too

OpenStudy (saifoo.khan):

Sorry havn't studied eco in my life.

OpenStudy (anonymous):

@amistre64

OpenStudy (mertsj):

"To solve for the equilibrium value of Y, one has to set Y equal to total expenditures, Z, and solve for Y." Don't know if that helps or not.

OpenStudy (anonymous):

is -7,437.5 the final answer?

OpenStudy (mertsj):

I don't know.

OpenStudy (anonymous):

so nobody knows how to do any of this? -_-

OpenStudy (mertsj):

I got 29750 but i know nothing about it.

OpenStudy (mertsj):

Y=2500+.8Y+2225+2000-775 and I solved it for Y. Have no idea if it makes any sense or not.

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

i added everything that can be added and got 0.80Y+5950

OpenStudy (anonymous):

then i set it equal to zero, subtracted, and then divided

OpenStudy (anonymous):

@TuringTest

OpenStudy (mertsj):

.2Y=5950 Y=29750

OpenStudy (mertsj):

2500+2225+2000-775=5950

OpenStudy (anonymous):

its .80Y

OpenStudy (mertsj):

1Y-.8Y=.2Y

OpenStudy (mertsj):

.8Y is the same as .80Y

OpenStudy (anonymous):

where are you getting this 1Y from?

OpenStudy (mertsj):

\[\frac{8}{10}=\frac{80}{100}\]

OpenStudy (mertsj):

To solve for the equilibrium value of Y, one has to set Y equal to total expenditures, Z, and solve for Y."

OpenStudy (mertsj):

So I said Y= all that stuff I added up.

OpenStudy (anonymous):

isnt it supposed to be -29750?

OpenStudy (mertsj):

\[Y=2500+.80Y+2225+2000-775\] \[1Y=(2500+2225+2000-775)+.80Y\] \[1Y=5950+.80Y\] \[1.00Y-.80Y=5950\] \[.20Y=5950\] \[Y=\frac{5950}{.20}\] \[Y=29750\]

OpenStudy (anonymous):

help with #10 in attachment above ^^^^^ ?

OpenStudy (mertsj):

Was that one right?

OpenStudy (mertsj):

Aggregate expenditure = consumption expenditure + Investment expenditure + Government expenditure+Exports-Imports.

OpenStudy (anonymous):

I dont know if it's right. It's not online hw. It's written hw.

OpenStudy (anonymous):

so aggregate expenditure would be 29,750 for both and the unplanned change in inventories would be 10,750 and 18,250?

OpenStudy (anonymous):

@mertsj

OpenStudy (mertsj):

I thought the 29750 was the equilibrium GDP.

OpenStudy (anonymous):

the formulas for both look the same

OpenStudy (anonymous):

i dont know. im confused.

OpenStudy (mertsj):

what is NX in your example?

OpenStudy (anonymous):

its all in the attachment above

OpenStudy (mertsj):

What does NX stand for?

OpenStudy (anonymous):

net exports?

OpenStudy (mertsj):

If that is true, why don't you just use the equation for Aggregate Expenditures to find Aggregate Expenditures? Aggregate expenditure = consumption expenditure + Investment expenditure + Government expenditure+Exports-Imports.

OpenStudy (anonymous):

that is what i did, no?

OpenStudy (anonymous):

o and the numbers i got above for the unplanned change in inventories should be -10,750 and -18,250... both negatives

OpenStudy (mertsj):

Does it turn out to be the same as the Equilibrium GDP?

OpenStudy (mertsj):

Seems like that would be reasonable.

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

isn't there any way to find out if you are right or wrong?

OpenStudy (anonymous):

not really. unless someone else knows how to do all of this.

OpenStudy (mertsj):

So what are you going to do with these numbers after you find them?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

??? Its a hw assignment that i have to turn in.

OpenStudy (anonymous):

you still there?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

I have been looking on the internet and all the equations say that Consumption is equal to some decimal times disposable income + some constant. So I wish you would look in your book or whatever you study from and find out exactly what that equation C=2500 +.80Y means. What does Y represent?

OpenStudy (mertsj):

What is the name and author of your book?

OpenStudy (anonymous):

R. Glenn Hubbard Anthony Patrick O'Brien Macroeconomics 3rd Edition Pearson

OpenStudy (anonymous):

Y stands for GDP I think

OpenStudy (anonymous):

Gross Domestic Product

OpenStudy (mertsj):

Better look it up and find out for sure. They are different types of GDP

OpenStudy (anonymous):

Just says GDP

OpenStudy (mertsj):

Then look up Equilibrium GDP and see what it says about that. there is probably an example in your book exactly like the one in this example.

OpenStudy (anonymous):

It says Aggregate Expenditure = GDP

OpenStudy (mertsj):

What does it say about Equilibrium GDP and Potential GDP?

OpenStudy (mertsj):

It asks for a name and password.

OpenStudy (anonymous):

IF … THEN … AND … Aggregate expenditure is equal to GDP inventories are unchanged the economy is in macroeconomic equilibrium. Aggregate expenditure is less than GDP inventories rise GDP and employment decrease. Aggregate expenditure is greater than GDP inventories fall GDP and employment increase.

OpenStudy (anonymous):

This is the relationship between aggregate expenditures and GDP

OpenStudy (mertsj):

So AE is equal to Equilibrium GDP. Would you agree?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So for number 19 plug in 19000 for Y in the AE equation. Did you do that/

OpenStudy (mertsj):

Then I don't know about the unplanned change in inventory.

OpenStudy (anonymous):

the unplanned change in inventory is GDP - Aggregate Expenditure

OpenStudy (mertsj):

Oh. I see. It is REAL GDP -AE

OpenStudy (mertsj):

So calculate AE and subtract from 19000

OpenStudy (anonymous):

I did and got -10,750 and -18,250

OpenStudy (mertsj):

So you're done?

OpenStudy (anonymous):

Yeah. I just wasn't sure I was doing everything correctly.

OpenStudy (mertsj):

I'm checking those Unplanned change in inventory figures.

OpenStudy (mertsj):

I got -2150 for the first one.

OpenStudy (anonymous):

show me what you are calculating on here

OpenStudy (mertsj):

AE=2500+.8(19000)+2225+2000-775

OpenStudy (mertsj):

Is that what you think?

OpenStudy (mertsj):

AE=2500+.8(19000)+2225+2000-775 = 21150 And 19000-21150 = -2150

OpenStudy (anonymous):

thats not how i did it... i used the 29750 i got for GDP (because GDP=Expenditures) and then i used the formula GDP-Expenditures to get the unplanned change in inventory... so: new GDP (19,000 and 11,500) - (29,750 for both)

OpenStudy (anonymous):

Getting: -10,750 and -18,250

OpenStudy (mertsj):

Oh. Ok. I'm sure you understand it much better than I do. I interpreted the problem differently.